Please help me about this problem. Roni is 30 years old and Grace is 22 years old. How many years ago was Roni twice as old as Grace?
Answer by ikleyn(52788)

.

The difference in their ages is, was and will be 30-22 = 8 years.


Hence, Roni was twice as old as Grace, when Grace was 8 years old and Roni was 16 years old.


It was 22-8 = 14 years ago.


Roni was 30-14 = 16 years old then.

Solved, answered and explained.

---------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.